# Interpretable LLM-based Table Question Answering

### v2

@@ 3.1. Planning in Natural Language
3.2. Converting Step to SQL
We harness the versatile capabilities of LLMs as Text-to-SQL converters (Hong et al., 2024) to translate each atomic step into its corresponding SQL command. There are two main reasons for us to use Step-to-SQLFigure 2). First, executing steps through SQL commands significantly improves accuracy compared to having LLMs perform the steps directly (see Section 4.2.3). Second, using SQL makes table transformations trackable, allowing us to identify which specific table cells are utilized during the decision-making process. In Figure 2, Step-to-SQL takes the current intermediate table and an atomic step as input and generates an SQL command. We use the following prompt to guide LLMs in converting text to SQL queries.:
@@ 3.2. Converting Step to SQL
3.3. Executing Plan of SQL Commands
Once each step is translated into SQL, we execute the plan using an SQL engine (sqlite3(Gaffney et al., 2022). The execution proceeds sequentially: the output table of an SQL command becomes the input of the next, chaining the transformations specified by the generated plan (Figure 2234) into model reasoning process.. In contrast to end-to-end and hybrid QA methods that rely on the black-box reasoning of LLMs to generate the final answer in Figure 1a & c, POS maintains the interpretability throughout. We provide details of the prompt engineering techniques for POS in Appendix G.
3.4. Generating Explanations for Table QA
@@
Despite the growing body of literature on Table QA and the pressing need for interpretability, it is surprising that no formal approach has been established to generate explanations for LLM-based Table QA systems. Previous studies have largely focused on improving accuracy, often overlooking the need to explain the decision-making processes of Table QA models (Wu and Feng, 2024; Kong et al., 2024). To address this gap, we propose an approach to generate explanations for Table QA using attribution maps–the main medium for explaining AI decisions to humans in various domains, including image classification (Colin et al., 2022), text analysis (Hase and Bansal, 2020), or time series (Theissler et al., 2022).
Per step, we generate an attribution map (Montavon et al., 2018; Nguyen et al., 2021a) for the current intermediate table, highlighting the specific information selected to progress to the next step (see Figure 2). The row and column selections are yellow-highlighted, while cells that match the specific condition in the step are green-highlighted. Finally, we combine the three components: (1) intermediate tables; (2) attribution maps; and (3) step descriptions to create an explanation shown to users to explain, for example, why the answer is True in Figure 2. Our explanation can be viewed as chain of attribution maps, helping users visually follow the sequence of transformations and understand how each table cell contributes to the final answer. Our visualization approach is detailed in Section A.1.
@@
Table 1. (a) Preference Rankings for XAI methods given by LLM judges on TabFact. Lower values indicate better rankings $\downarrow$ (1 = best, 4 = worst). (b) Simulation Accuracy $\uparrow$ (%) of LLM and human judges for on TabFact.
@@
Table 2. Verification Accuracy (%) $\uparrow$ of LLM judges on TabFact and WikiTQ.
@@
4. Experiments
We conduct experiments using three popular and standard Table QA benchmarks: TabFact (Chen et al., 2020), WikiTQ (Pasupat and Liang, 2015), and FeTaQA (Nan et al., 2022).
TabFact is a fact verification dataset in which each statement associated with a table is labeled TRUE or FALSE. We use the cleaned TabFact dataset from Wang et al. (2024) and evaluate Table QA methods with binary classification accuracy on the 2,024-sample test-small set.
WikiTQ is a question-answering dataset where the goal is to answer human-written questions using an input table. Using the dataset and evaluation scripts from Ye et al. (2023), we assess model denotation accuracy (whether the predicted answer is equal to the ground-truth answer) on the 4,344-sample standard test set.
FeTaQA is a free-form Table QA dataset where the task is to generate free-form natural language responses based on information retrieved or inferred from a table. We evaluate models on the 2,003-sample standard test set using BLEU and ROUGE.
4.1. Evaluating Explanations for Table QA
Baselines.
We select Text-to-SQL (Rajkumar et al., 2022), DATER (Ye et al., 2023), and Chain-of-Table (CoTable) (Wang et al., 2024) to generate explanations. These methods are selected for their high accuracy, interpretability, and reproducibility (see Appendix A for more details). Later, we compare POS with self-explanations (Madsen et al., 2024), where LLMs are prompted to provide explanations for their own answers.
Evaluation measures.
Following the two evaluation settings proposed by Doshi-Velez and Kim (2017), we evaluate explanations using both qualitative and quantitative measures. These evaluations are conducted by human users as well as LLM judges (LLM-as-XAI-judge).
In qualitative evaluation, XAI judges (humans and LLMs) receive the input, output, and explanation. They are then asked to rank the explanations based on perceived quality, referred to as Preference (Ramaswamy et al., 2023; Yang et al., 2024). This setup directly compares explanations with rubrics: clarity, coherence, and helpfulness in enabling XAI judges to understand the model’s reasoning (see Appendix F).
Final rankings for each method are computed as the average preference ranking across 707 TabFact samples. Only samples where all methods generate either correct or incorrect answers are included, ensuring that the rankings are not biased toward methods with higher TabFact accuracy. Lower ranking values indicate better quality $\downarrow$ (1 = best, 4 = worst, since we test four explanation methods).
In quantitative evaluation, XAI judges receive input and explanation and are tasked with evaluating two measures of understanding: model decision boundary (via forward simulation) and model error (via model prediction verification) (Chen et al., 2023a).
In the first task, XAI judges predict the model’s behaviors (i.e., its decision boundary) based on the explanations, regardless of the ground truth. This task is known as Forward Simulation (Doshi-Velez and Kim, 2017; Hase and Bansal, 2020; Chen et al., 2022; Mills et al., 2023) or Simulation for short.
In the second task, we additionally present the model output to the XAI judges and ask them to analyze the explanation to verify whether the AI model’s output is correct or incorrect. This task is known as Model Prediction Verification (Nguyen et al., 2021b; Taesiri et al., 2022) or Verification for short.
4.1.1. Evaluating explanations with humans
Motivation.
Human evaluation is considered the gold standard for evaluating AI explanations, as humans are the ultimate users who work with AI models (Doshi-Velez and Kim, 2017). We aim to study how explanations help humans in understanding then predicting model behaviors via Forward Simulation.
Participants and Data.
We recruit 32 volunteers, all of whom are undergraduate, master, or Ph.D. students in Computer Science. In each session, users select one of four explanation methods and complete 10 samples, with the option to participate in multiple sessions. In the end, we collect 800 responses ($\approx 200$ per explanation method). We show our human study interface in  Appendix J.
Table 3. LLM–human agreement (%) on forward simulation for TabFact. Each cell indicates how often an LLM’s decisions align with the human decisions on the same subset of samples (listed in Human Samples). Higher values suggest stronger agreement between LLMs and humans.
Method GPT4o-mini Qwen2.5-72B-Inst Llama-3.1-405B-Inst Human Samples
Text-to-SQL 71.24% 79.08% 75.16% 153
DATER 81.98% 83.14% 76.74% 172
CoTable 79.10% 82.49% 80.79% 177
POS 88.30% 90.59% 89.41% 171
4.1.2. Evaluating explanations with LLMs
Motivation.
The use of LLMs trained to align with human preference (Ouyang et al., 2022) as judges has been gaining attention due to their strong correlation with human judgments (Dubois et al., 2024; Zheng et al., 2023; Liu et al., 2023; Mills et al., 2023; Fernández-Becerra et al., 2024; Poché et al., 2025). This evidence makes LLMs a promising, scalable solution for evaluating explanations, particularly in tasks like Table QA, where the information is still text-based but structured. We are motivated to leverage LLM judges for all three tasks: Preference Ranking, Forward Simulation, and Model Prediction Verification.
LLM judges.
Inspired by previous works showing the effectiveness of OpenAI’s GPT models as reliable judges (Zheng et al., 2023; Liu et al., 2023; Dubois et al., 2024), we utilize 3 OpenAI’s models: gpt-4-turbo-2024-04-09, gpt-4o, and gpt-4o-mini to judge Table QA explanations. Additionally, we present an evaluation of open-source LLMs in Section 4.1.5 and the detailed setup of the LLM-as-XAI-Judge experiments in Appendix F.
4.1.3. Findings of explanation evaluation
@@
In Table 1(a), POS explanations consistently receive the best ranks from all LLM judges. Specifically, our proposed explanations achieve average ranks of 1.33, 1.55, and 1.01 from GPT-4, GPT-4o-mini, and GPT-4o respectively, substantially outperforming CoTable, DATER, and Text-to-SQL.
This shows that POS explanations are regarded by LLM judges as the most clear, coherent, and helpful for understanding the model’s reasoning process (see rubrics in Appendix F). The preference for POS highlights its potential to gain trust, a key factor for AI adoption in real-world applications (Doshi-Velez and Kim, 2017; Lipton, 2018).
POS is most effective for understanding model behaviors.
Table 1(b) shows that POS effectively helps human and LLM judges predict the model behaviors. Specifically, human judges achieve 93.00% with POS explanations, outperforming other methods such as DATER (86.50%) and CoTable (84.29%). Similarly, across all LLM judges, POS consistently yields the highest accuracy, with improvements ranging from 5% $\rightarrow$ 6% over the next best XAI method.
@@
We have shown that POS explanations are more informative and predictive than other baselines via forward simulation. Building on these findings, we extend our evaluation to a real-world scenario—identifying whether AI outputs are correct or incorrect, an inherent daily task for AI users (Alufaisan et al., 2021; Nguyen et al., 2021b). We perform experiments on TabFact and WikiTQ, comparing the verification accuracy with different XAI methods in  Table 2.
We find that POS is the best method in 5/6 evaluation settings. In addition, the improvements between DATER and POS are more profound in TabFact compared to WikiTQ, suggesting that the utility of explanations is influenced by the nature of the task. As a more complex dataset, WikiTQ makes it inherently more difficult to verify the predictions using explanations alone. In Section 4.1.6, we present an ablation study that further studies the factors driving POS’s improved interpretability.
@@
Using Table 1Table 2, we perform a correlation analysis to study whether qualitative rankings inform quantitative measurements in TabFact. Since lower rankings in  Table 1(a) indicate better explanations, we invert the rankings to align higher simulation/verification accuracy with better preference.
@@
We find statistically significant positive Pearson correlations between preference rankings vs. simulation accuracy ($r=0.7865$, $p=0.0024$) and vs. verification accuracy ($r=0.7035$, $p=0.0107$). These findings suggest that in Table QA, our proposed rubrics based on perceived quality (see Appendix F) can effectively filter high-quality explanations. This allows for efficient pre-selection of explanations to present to users, reducing the need for expensive and time-consuming user studies.
Comparison with post-hoc self-explanations
@@
Finally, we investigate whether a simple, post-hoc self-explanation approach could rival POS in interpretability. Table 10 compares POS with a widely used “self-explanation” method, in which the LLM first generates an answer and then retrospectively explains its reasoning ( Figure 8).
@@
We find that self-explanation remains substantially less interpretable than POS, largely due to the lack of faithfulness in post-hoc explanations (Madsen et al., 2024; Chen et al., 2023b; Agarwal et al., 2024). By contrast, POS grounds each explanation in offline transformation steps (via SQL execution), yielding absolute faithfulness and comprehensiveness, which are especially valuable for model simulation or verification. Please refer to Appendix B for details.
4.1.4. LLM–Human agreement in XAI evaluation
In Table 3, we report the instance-level agreement between LLM-based and human forward simulations on the same samples for four XAI methods. This metric differs from accuracy: it specifically measures how often LLM judges and humans arrive at the same decisions, given identical explanations, even if the underlying model’s final predictions might or might not be correct. We find that:
  • POS consistently yields the highest LLM–human agreement across all tested language models, reaching up to 90.59%. This suggests that when explanations are faithful and grounded—e.g., via atomic SQL steps—both LLMs and humans converge on similar decisions in evaluating explanations.
  • CoTable, DATER, and Text-to-SQL show lower agreement (70–83%), suggesting that they are either less informative or faithful explanations.
The “Human Samples” column indicates that each method was evaluated by at least 150 human annotations. Overall, our findings suggest the idea that high-quality, step-by-step explanations can achieve consensus between humans and LLM evaluators. In practice, this means LLMs may serve as scalable proxies for early-phase evaluation of Table QA explanations, reducing the burden of comprehensive user studies while still approximating human judgments.
4.1.5. Evaluating explanations with open-source LLMs
@@
So far, our evaluation relies primarily on GPT-family models. This can restrict the generalizability of our findings in  Section 4.1.3. To address this, we introduce two additional model families—Qwen and Llama—to confirm whether the observed trends in explanation quality persist across multiple LLM families.
Setup.
@@
We use TabFact as the benchmark dataset, focusing on two tasks: Forward Simulation and Model Prediction Verification. In both tasks, we test four methods—Text-to-SQL, DATER, CoTable, and POS —under open-source Qwen2.5-72B-Inst and Llama-3.1-405B-Inst from SambaNova. [^1] https://sambanova.ai/
@@
Table 4. Evaluating Table QA explanations (%) on TabFact with open-source models.
@@
Results.
We report the evaluation results on open-source LLMs in Table 4. Regarding verification accuracy, POS consistently delivers top performance, while the Text-to-SQL approach remains the weakest across all model families. In forward simulation, POS outperforms other baselines on Llama, although DATER slightly exceeds POS for Qwen2.5 (91.38% vs. 90.59%). In general, we found that the interpretability provided by POS generalizes well to open-source LLMs.
4.1.6. Ablation study on POS interpretability
To better understand which components of POS most strongly contribute to its interpretability, we perform an ablation study on two datasets: TabFact and WikiTQ. In particular, we focus on the model prediction verification task in which an XAI judge (here, GPT-4o-mini) is shown POS’s explanations and asked to verify whether the final answer is correct or incorrect.
Setup.
We remove (i.e., ablate) one of the following three core components in POS and observe the resultant drop (or change) in verification accuracy:
  • Atomic Operations: Instead of enforcing single-condition transformations, we allow complex multi-condition steps.
  • NL Planning: We replace the natural-language planning with a direct prompt that asks the LLM to generate a sequence of SQL commands to solve the query.
  • SQL Execution: We replace the SQL-based transformations with direct black-box LLM-based transformations.
Results.
Table 5 shows the verification accuracy (%) on TabFact and WikiTQ in each ablation setting. We see that removing SQL execution leads to the biggest interpretability drop (down to 64.19% on TabFact), while removing the plan also yields a substantial decrease. Atomic operations, though beneficial, exhibit the smallest impact when removed.
Table 5. Verification Accuracy (%) on TabFact and WikiTQ. Removing each component from POS reveals its impact on interpretability: SQL execution proves most critical, followed by the plan, and then atomic operations.

### v3

@@ 3.1 Planning in Natural Language
3.2 Converting Step to SQL
Once we have the plan in natural language, the next stage is to translate each step into an executable SQL query. We leverage the LLMs’ capability as a Text-to-SQL converter (Hong et al., 2024) to perform this conversion for each step. There are two main reasons to use a dedicated Step-to-SQL module (Fig. 2). First, executing the reasoning steps via SQL commands greatly improves the correctness of each step compared to letting LLMs manipulate the table implicitly in its hidden state (see our analysis in Appendix C). In particular, offloading the computation to an SQL query ensures each operation is carried out exactly and eliminates errors from the LLMs’ reasoning. Second, using SQL makes the sequence of transformations fully trackable: for every step, we know precisely which table entries were accessed or modified, since the SQL query specifies those conditions. This allows us to identify what information each step used.
In Fig. 2, the Step-to-SQL module takes the current intermediate table (initially the original table $T$) and one atomic step as input, and outputs the corresponding SQL query. We steer an LLM to generate an SQL with a specialized prompt template, as illustrated below:
@@ 3.2 Converting Step to SQL
After converting all steps, POS executes them sequentially using an SQL engine (we use SQLite3 (Gaffney et al., 2022) in our implementation). The intermediate result of each SQL command becomes the input for the next step, faithfully carrying out the transformations specified by the plan (see the chain Fig. 2 2$\rightarrow$ 3$\rightarrow$ 4). In contrast to end-to-end or “chain-of-thought” Table QA methods that rely on an LLM’s latent reasoning to jump to the final answer (as in Fig. 1a, c), POS maintains interpretability throughout the process: the answer is obtained through another SQL operation whose effects can be understood.
3.3 Generating Explanations for Table QA
@@
Despite the need for interpretability, prior work has not established a formal method to generate human-understandable explanations for LLM-based Table QA models. Most studies focus on improving accuracy and leave the model’s decision-making process opaque (Wu & Feng, 2024; Kong et al., 2024; Cheng et al., 2025). To bridge this gap, we propose an approach to generate explanations for Table QA using attribution maps—the main medium for explaining AI decisions to humans in various domains, including image classification (Colin et al., 2022), text analysis (Hase & Bansal, 2020), and time series (Theissler et al., 2022).
For each step, we create an attribution map on the intermediate table that highlights the information used in that step. Specifically, as we apply a transformation, we mark the rows and columns that were selected or affected in yellow, and we highlight the cells that satisfy the step’s condition in green. By doing this for every step (see Fig. 2), we obtain intermediate tables annotated with highlights indicating which data contributed to the final answer. Finally, we present the explanation as a chain of attribution maps: each step is shown alongside its highlighted intermediate table. This allows a user to visually follow the process of reasoning. In practice, the explanation shows which cells were used when, making it clear why, for example, the final answer is True in the running example of Fig. 2. We also apply this visualization approach to other Table QA methods, such as CoTable (Wang et al., 2024) and DATER (Ye et al., 2023), with additional examples provided in Appendix A.
4 Experiments
We conduct experiments using three popular and standard Table QA benchmarks: TabFact (Chen et al., 2020), WikiTQ (Pasupat & Liang, 2015), and FeTaQA Nan et al. (2022). TabFact is a fact verification dataset in which each statement associated with a table is labeled TRUE or FALSE. We use the cleaned TabFact dataset from Wang et al. (2024) and evaluate Table QA methods with binary classification accuracy on the 2,024-sample test-small set. WikiTQ is a question-answering dataset where the goal is to answer human-written questions using an input table. Using the dataset and evaluation scripts from Ye et al. (2023), we assess model denotation accuracy (whether the predicted answer is equal to the ground-truth answer) on the 4,344-sample standard test set. FeTaQA is a free-form Table QA dataset where the task is to generate free-form natural language responses based on information retrieved or inferred from a table. We evaluate models on the 2,003-sample standard test set using BLEU and ROUGE.
4.1 Evaluating Explanations in Table QA
Baselines.
We select Text-to-SQL (Rajkumar et al., 2022), DATER (Ye et al., 2023), and Chain-of-Table (CoTable) (Wang et al., 2024) to benchmark the interpretability of POS. They are chosen for their high performance, interpretability, and reproducibility (see details of baseline methods in Appendix A). Later, we compare POS with self-explanations (Madsen et al., 2024), where LLMs are prompted to provide explanations for their own answers.
Evaluation measures.
To evaluate the quality of explanations, we involve both human judgments and LLM-based judgments (which we denote as LLM-as-XAI-judge). We follow the two complementary evaluation perspectives (qualitative and quantitative) proposed by Doshi-Velez & Kim (2017).
In the qualitative evaluation, the judge is shown the table, the question, the model’s answer, and the model’s explanation. The task is to rank the explanations from different methods by overall perceived preference (to identify which explanations users are most likely to use in practice). This preference rating is based on a clearly defined rubric comprising three criteria: clarity, coherence, and usefulness in understanding the model’s reasoning. We refer to this as a Preference Ranking task, following prior works on comparative explanation evaluation (Ramaswamy et al., 2023; Yang et al., 2024; Zhang et al., 2025; Wazzan et al., 2025). For each table question, we collect a ranking of the four methods’ explanations (1 = best, 4 = worst). To ensure fairness, we only consider test samples where all methods being compared got the question either correct or incorrect (so that judges are comparing explanations for answers of the same correctness). We aggregate rankings over 707 TabFact questions meeting this criterion, computing the average rank for each method (lower is better $\downarrow$).
In the quantitative evaluation, the judge is asked to perform tasks that objectively measure how well the explanation informs them about the model’s behaviors. We use two standard tasks for evaluating explanations: Forward Simulation (Doshi-Velez & Kim, 2017; Hase & Bansal, 2020; Chen et al., 2022; Mills et al., 2023) and Model Prediction Verification (Nguyen et al., 2021; Taesiri et al., 2022; Chen et al., 2023a).
In Forward Simulation, the judge is given the table, question, and explanation without the model’s answer, and must predict what answer the model’s output would be. This evaluates how clearly the explanation communicates the model’s decision boundary or reasoning process to users.
In Prediction Verification, the judge is given the table, question, model’s answer, and explanation, then must decide whether the model’s answer is correct or not based on the explanation. This measures how well the explanation justifies the model’s prediction (e.g., can the judge catch model errors using explanations?).
We compute performance for these tasks as the percentage of samples in which the judge makes a correct decision, reported as Simulation Accuracy and Verification Accuracy, respectively.
4.1.1 Evaluating explanations with human users
Motivation.
Human evaluation is considered the gold standard for evaluating AI explanations, as humans are the ultimate users who work with AI models (Doshi-Velez & Kim, 2017). We aim to study how explanations help humans in understanding then predicting model behaviors via Forward Simulation.
Participants and Data.
We recruit 32 volunteers, all of whom are undergraduate, master, or Ph.D. students in Computer Science. In each session, users select one of four explanation methods and complete 10 samples, with an option to participate in multiple sessions. We collect 800 responses ($\approx 200$ per method).
4.1.2 Evaluating explanations with LLMs
Motivation.
The use of LLMs trained to align with human preference (Ouyang et al., 2022) as judges has been gaining attention due to their strong correlation with human judgments (Dubois et al., 2024; Zheng et al., 2023; Liu et al., 2023; Mills et al., 2023; Fernández-Becerra et al., 2024; Poché et al., 2025). This makes LLM judge a promising, scalable solution for evaluating explanations, particularly in tasks like Table QA, where the information is still text-based yet structured. Thus, we are motivated to leverage LLM judges for all three tasks: Preference Ranking, Forward Simulation, and Model Prediction Verification.
LLM judge.
Inspired by recent works showing the effectiveness of OpenAI’s instruction-tuned GPT models as reliable judges (Zheng et al., 2023; Liu et al., 2023; Dubois et al., 2024), we utilize 3 OpenAI’s LLMs: gpt-4-turbo-2024-04-09, gpt-4o, and gpt-4o-mini to evaluate Table QA explanations. We ensure the prompts encourage the model to follow the given instructions strictly (see Appendix G for the exact prompt templates and calibration procedures). We also explore the use of open-source LLMs as judges in Sec. 4.1.4.
@@ 4.1.2 Evaluating explanations with LLMs
Table 1: (a) Preference Rankings for explanation methods given by LLM judges on TabFact. Lower values indicate better rankings $\downarrow$. (b) Simulation Accuracy $\uparrow$ (%) of LLM and human judges on TabFact.
@@
Table 2: Verification Accuracy $\uparrow$ (%) of LLM judges on TabFact and WikiTQ.
@@
4.1.3 Findings from the evaluation of explanations
@@
In Tab. 1(a), POS explanations consistently receive the best ranks from all LLM judges. Specifically, our explanations achieve average ranks of 1.33, 1.55, and 1.01 from GPT-4, GPT-4o-mini, and GPT-4o; respectively, substantially outperforming CoTable, DATER, and Text-to-SQL.
This shows that POS explanations are regarded by LLM judges as the clearest, most coherent, and most helpful for understanding the Table QA model’s reasoning process (see the rubrics in Appendix G). In practice, this preference could translate into increased trust—a key factor for AI adoption in high-stakes domains (Doshi-Velez & Kim, 2017).
POS is most effective for predicting model behaviors.
Tab. 1(b) shows that POS effectively helps human and LLM judges predict the model behaviors. Specifically, human judges achieve 93.00% with POS explanations, outperforming other methods such as DATER (86.50%) and CoTable (84.29%). Similarly, across all LLM judges, POS consistently yields the highest accuracy, with improvements ranging from 5% $\rightarrow$ 6% over the second best methods.
@@
We perform this experiment on TabFact and WikiTQ, comparing the verification accuracy with different explanation methods in  Tab. 2. We find that POS is the best method in five out of six settings. In addition, the improvements between DATER and POS are more pronounced in TabFact compared to WikiTQ, suggesting that the effectiveness of explanations is influenced by the nature of the task. As a more complex dataset, WikiTQ makes it inherently more difficult for the judges to verify the predictions. Please note that we do not compare POS with Text-to-SQL and CoTable on WikiTQ, as these methods have not been publicly released for this benchmark.
Later in Sec. 4.1.6, we present an ablation study that further studies the factors driving POS’s improved interpretability.
@@
Using Tab. 1Tab. 2, we perform a correlation analysis to study whether qualitative preference rankings inform quantitative measures. Since lower rankings in  Tab. 1(a) indicate higher-quality explanations, we invert the rankings to align higher simulation/verification accuracy with higher preference.
@@
Interestingly, we find statistically significant positive Pearson correlations between preference rankings vs. simulation accuracy ($r=0.7865$, $p=0.0024$) and vs. verification accuracy ($r=0.7035$, $p=0.0107$). These high correlations suggest that in Table QA, our proposed rubrics based on perceived quality (see Appendix G) can effectively identify high-utility explanations. This allows for efficient pre-selection of explanations to present to users, significantly reducing the need for expensive and time-consuming user studies.
Comparison with post-hoc self-explanations.
@@
Finally, we investigate whether a simple, post-hoc self-explaining approach could rival POS in interpretability. Tab. 10 compares POS with a widely used “self-explanation” method, in which the LLM first generates an answer and then retrospectively explains its answer (see an example in Fig. 8).
@@
We find that self-explanation remains substantially less effective than POS in all three interpretability benchmarks, largely due to the lack of faithfulness in post-hoc explanations (Madsen et al., 2024; Chen et al., 2023b; Agarwal et al., 2024). By contrast, POS grounds its explanations in offline transformation steps (via SQL execution), yielding absolute faithfulness and comprehensiveness, which are especially valuable for model simulation or verification. We provide more details in Sec. A.3.
4.1.4 Evaluating explanations with open-source LLMs
@@
So far, our evaluation relies on GPT-family models. This can restrict the generalizability of our findings in  Sec. 4.1.3. To address this, we use two additional model families—Qwen and Llama—to confirm whether the observed trends in explanation quality persist across multiple LLM families.
@@
We use TabFact as the benchmark dataset, focusing on two tasks: Forward Simulation and Model Prediction Verification. In both tasks, we test four methods—Text-to-SQL, DATER, CoTable, and POS —under open-source Qwen2.5-72B-Inst and Llama-3.1-405B-Inst hosted by SambaNova [^1] https://sambanova.ai/.
@@
Table 3: Evaluating Table QA explanations (%) on TabFact with open-source models.
@@
We report the evaluation results on open-source LLMs in Tab. 3. In Verification, POS consistently delivers top performance, while Text-to-SQL remains the worst across all model families. In Forward Simulation, POS outperforms other baselines on Llama, although DATER slightly exceeds POS with Qwen2.5 (91.38% vs. 90.59%). In general, we find that the interpretability provided by POS translates well to open-source LLMs.
4.1.5 LLM–Human agreement in XAI evaluation
In Tab. 4, we report the instance-level agreement between LLM-based and human forward simulations on the same samples for four explanation methods. This metric differs from accuracy; it measures how often LLM judges and humans arrive at the same decision when given identical information (i.e., the input and an explanation). We find that: First, POS consistently yields the highest LLM–human agreement across all tested LLM judges, reaching up to 90.59%. This suggests that when explanations are faithful and grounded—e.g., via atomic SQL steps—both LLMs and humans converge on similar decisions in evaluating explanations. Second, baseline methods exhibit lower agreement (71–83%), likely because their explanations are less informative or less faithful, leading LLM judges and humans to follow different reasoning paths. Practically, these high agreements (71–90%) suggest that LLMs may serve as effective proxies for early-phase evaluation of Table QA explanations, reducing the high cost of user studies while still approximating human judgments.
Table 4: LLM–human agreement (%) on forward simulation for TabFact. Each cell indicates how often an LLM’s decisions align with the human decisions on the same subset of samples (listed in Human Samples). Higher values suggest stronger agreement between LLMs and humans.
Method GPT-4o-mini Qwen2.5-72B-Inst Llama-3.1-405B-Inst Human Samples
Text-to-SQL 71.24% 79.08% 75.16% 153
DATER 81.98% 83.14% 76.74% 172
CoTable 79.10% 82.49% 80.79% 177
POS 88.30% 90.59% 89.41% 171
4.1.6 Ablation study on POS interpretability
To better understand which components of POS (Sec. 3) most strongly contribute to its interpretability, we perform an ablation study on both TabFact and WikiTQ. We focus on the model prediction verification in which an XAI judge (here, GPT-4o-mini) is shown POS’s explanations and asked to verify the final answer.
We remove (i.e., ablate out) one of the following three core components in POS and observe the resultant drop (or change) in verification accuracy: (i) Atomic operations: Instead of enforcing single-condition transformations, we allow complex multi-condition steps. (ii) NL planning: We replace the natural-language planning with a direct prompt that asks the LLM to generate a sequence of SQL commands to solve the question. (iii) SQL execution: We replace the SQL-based transformations with direct black-box LLM-based transformations.
Tab. 5 shows the verification accuracy in each ablation setting. We find that removing SQL execution leads to the biggest interpretability drop (from 76.74 to 64.19% on TabFact), while removing the natural-language planning also leads to a substantial decrease. Atomic operations, though beneficial, exhibit the smallest impact when removed.
Table 5: Verification Accuracy (%) on TabFact and WikiTQ. Removing each component from POS reveals its impact on interpretability: SQL execution is most critical, followed by the plan, and then atomic operations.